OpenStudy (anonymous):

help please..A 18.01 g sample of N2 reacts with 3.02g of H2 to form ammonia(NH3). what mass of ammonia is formed?

OpenStudy (anonymous):

I have the balanced equation N2+3H2->2NH3 what should i do next?

OpenStudy (chmvijay):

u should find limiting reagent for the reaction N2 or H2 ?

OpenStudy (anonymous):

how would I find it?

OpenStudy (chmvijay):

see the balanced equation for one mole of N2 u need 3 mole of H2 right

OpenStudy (anonymous):

yes

OpenStudy (chmvijay):

H2 and N2 molar mass? convert the gram into moles present the question

OpenStudy (anonymous):

okay and if i find the mole of H2 and then multiply it by the ratio?

OpenStudy (chmvijay):

nope u find it first

OpenStudy (anonymous):

okay I got 0.64mole for N2 and 1.50 mole for H2. what do I do next?

OpenStudy (chmvijay):

good

OpenStudy (chmvijay):

so according to balanced equation u need 3 mole of H2 for 1 mole of N2 1.50 moles of H2 u need =1.5*1/3 =0.5 moles of N2

OpenStudy (chmvijay):

but u have 0.64 moles of N2

OpenStudy (chmvijay):

mean u have excess of N2 than required for ur H2

OpenStudy (chmvijay):

so H2 is the limiting reagent

OpenStudy (anonymous):

okay after I know H2 is the limiting reagent what do I do next?

OpenStudy (chmvijay):

NH3 molar mass ?

OpenStudy (anonymous):

17.04 NH3

OpenStudy (chmvijay):

sorry here N2 is limiting reagent :P my bad

OpenStudy (anonymous):

its okay

OpenStudy (chmvijay):

14 gramof N@2 give rise to 2*17 gramof ammonia than 18 gramof N2 =2*17*18 /14 =

OpenStudy (anonymous):

43.8?

OpenStudy (chmvijay):

its simple maths i dont wanna say this at east

OpenStudy (chmvijay):

its simple maths i dont wanna say this at least